How many time constants will elapse before the power delivered by the battery drops to half of its maximum value in an RC circuit?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Power = CV^2 = Q * V Now, QV/2 = QV * e^-t/RC ⇒ ½ = e^–t/RC ⇒ t/RC = -In 0.5 ⇒ -(-0.69) = 0.69